OpenStudy (123ab456c):

What is the side length of a cube with a volume of 0.064 m^3? A. 0.000262 m B. 0.04 m C. 0.4 m D. 4 m

OpenStudy (123ab456c):

@satellite73

OpenStudy (notshh):

take the cubed root of .064

jhonyy9 (jhonyy9):

because the volume of a cube = side*side*side so side^3 make cuberoot(0,064)= ? 0,4*0,4*04 = 0,16*0,4=0,064 from this result the choice C. is right sure